OpenStudy (anonymous):

can someone check this please. Differentiate the function y = f(x) = x squared−x divided by x .

OpenStudy (turingtest):

2x

OpenStudy (anonymous):

\[f(x) ={ {x^2 -x} \over x} ?\]

OpenStudy (anonymous):

my final answer i got was -x^2 - 2x divided by x^2

hero (hero):

Oh, lol

hero (hero):

That has to be it

OpenStudy (turingtest):

if f(x)=(x^2-x)/x then f'(x)=1

OpenStudy (anonymous):

\[f(x) = x - 1 \] \[f'(x) = 1 \]

hero (hero):

Yes, correct f'(x) = 1

OpenStudy (turingtest):

if that's how the question was written

hero (hero):

Ishaan did it the best way

OpenStudy (turingtest):

otherwise without parentheses it would be 2x

hero (hero):

It's not 2x

OpenStudy (turingtest):

if he meant x^2-(x/x) it is. the question needs parentheses.

OpenStudy (anonymous):

problem = x^2 - x all divided by x then using the quotient rule = (x^2-x) (1) - (2x-1)(x) divided by x ^2. then x^2-x-2x^2 - x all divided by x squared. sorry if thats hard to follow but the question is on a mock exam without an answer key and has no parantheses.

OpenStudy (anonymous):

derivative of [x^2-x]/x is 1

hero (hero):